OpenStudy (anonymous):

Decide weather the percent is reasonable and explain why or why not. Rainfall in Oregon this year is reported to be 160% of the average

OpenStudy (tkhunny):

Unclear. Not enough information. We need information concerning central tendency. 99 100 101 -- Average is 100. Likelihood of 160? 0% 90 100 110 -- Average is 100. Likelihood of 160? 0% 00 100 200 -- Average is 100. Likelihood of 160? NOT 0%

OpenStudy (anonymous):

IDK that is all the information the question has

OpenStudy (anonymous):

I am just in 7th grade math so I think they are just looking for a pretty basic answer.

OpenStudy (tkhunny):

MOst basic answer? 60% is amazing! I don't believe it. *** Read the Rest with Caution *** That IS the answer in the absence of any other information. One COULD argue, that annual rainfall records have been around for more than 100 years, therefore, my 3-point examples are rather untoward. However, how much language have we heard lately like this, "100-year storm" or "500-year event"? A lot! Anyone who is paying attention to data and statistics in the real world, these days, knows we are now paying MUCH more attention to the tails of the distribution. Reading this paper: http://www.climate.com/assets/LandingPageDocs/rainfallstudy2007.pdf We see: The top ten cities in annual rainfall volatility average 194% more volatility than the annual U.S. average, 50 inches of annual fluctuation from the 60 inch top ten most volatile average vs. 26 inches from the 35 inch national average. 50 inches!!! Guess what cities are in the top ten for fluctuations? Eugene, OR, Olympia, WA, Salem, OR, and Portland, OR. An astounding 4/10 right around the area we're talking about, Oregon. So, 160% of the average in Montana? Astounding! Something must be wrong. 160% of the average in Oregon? It would be big news, but maybe! I can't rule it out.

OpenStudy (tkhunny):

From that same report, we have: 4. Eugene, OR 17.6 Average 6.27 Standard Deviation. 1.6 * 17.6 = 28.16 (28.16 - 17.6)/6.27 = 1.68 <== Not an unlikely outcome AT ALL!

OpenStudy (anonymous):

Thanks for all your information I appreciate it!!!

OpenStudy (tkhunny):

Just for the record: Havre, MT 1.3 Average .63 Standard Deviation 1.6 * 1.3 = 2.08 (2.08-1.3)/0.63 = 1.24 Well, I stand corrected by the data. It's pretty likely in Montana, too.